OpenStudy (anonymous):

Given f(x) and g(x) below, find: (a) f + g (b) f – g (c) f*g (d) f/g Please show all of your work. f(x) = 3 – 4x g(x) = -7x + 4

OpenStudy (anonymous):

given f(x)=3 – 4x, g(x) = -7x + 4 a) f + g =f(x)+ g(x) = 3 – 4x+ -7x + 4= -11x+7 (we added or subtracted liked terms) b) f - g =f(x)- g(x) = 3 – 4x-( -7x + 4) = 3 – 4x+7x - 4 = 3x-1 ans.... (we added or subtracted liked terms)

OpenStudy (anonymous):

(c) f*g =f(x)* g(x) = (3 – 4x)* (-7x + 4) =-21x+28x^2 +12+16x collecting terms to get =28x^2 -5x+12 ans

OpenStudy (anonymous):

(d) f/g =(3 – 4x)/(-7x + 4)

OpenStudy (anonymous):

you rock. Thank you so much Mark!